OpenStudy (anonymous):

solve equation -9=2+y

OpenStudy (anonymous):

could use some help on this one

OpenStudy (anonymous):

It isn't that complicated buddy, you just need to isolate Y, understand? E.g. x + 2 = 10 + 2 -2 from both sides x = 10 (: Understand?

OpenStudy (anonymous):

yea thank you

OpenStudy (whpalmer4):

You can do just about anything to the equation, so long as you do it to both sides. So, here you subtract 2 from both sides to leave y alone on one side.

OpenStudy (anonymous):

Yup, go try it bud, tell us the answer you got. :D